# AI Agentic Programming: A Survey of Techniques, Challenges, and Opportunities

### v1

@@ Abstract.
AI agentic programming is an emerging paradigm in which large language models (LLMs) autonomously plan, execute, and interact with external tools like compilers, debuggers, and version control systems to iteratively perform complex software development tasks. Unlike conventional code generation tools, agentic systems are capable of decomposing high-level goals, coordinating multi-step processes, and adapting their behavior based on intermediate feedback. These capabilities are transforming the software development practice. As this emerging field evolves rapidly, there is a need to define its scope, consolidate its technical foundations, and identify open research challenges. This survey provides a comprehensive and timely review of AI agentic programming. We introduce a taxonomy of agent behaviors and system architectures, and examine core techniques including planning, memory and context management, tool integration, and execution monitoring. We also analyze existing benchmarks and evaluation methodologies used to assess coding agent performance. Our study identifies several key challenges, including limitations in handling long context, a lack of persistent memory across tasks, and concerns around safety, alignment with user intent, and collaboration with human developers. We discuss emerging opportunities to improve the reliability, adaptability, and transparency of agentic systems. By synthesizing recent advances and outlining future directions, this survey aims to provide a foundation for research and development in building the next generation of intelligent and trustworthy AI coding agents.
@@ 1. Introduction
The paradigm of software development is changing rapidly with the rise of large language models (LLMs) (hou2024largelanguagemodelssoftware, ). These models enable artificial intelligence (AI) systems that not only generate code (TOSEM25, ) but also understand task requirements, interact with development tools, and iteratively refine their outputs (FSE24_Alshahwan, ; qian2024chatdevcommunicativeagentssoftware, ). Recent studies suggest that software developers now use LLMs routinely to assist in daily coding tasks (hou2024largelanguagemodelssoftware, ; DBLP:conf/fose-ws/FanGHLSYZ23, ; gong2025language, ). Unlike traditional code generation tools (TraditionalCodeGenration, ) that respond to a single prompt with a static code snippet, emerging AI coding agents are designed to operate within dynamic software environments, performing iterative, tool-augmented tasks to achieve complex goals.
@@
This shift has given rise to a new programming paradigm known as AI agentic programming, where LLM-based coding agents can autonomously plan, execute, and refine software development tasks (gao2023retrieval, ; novikov2025alphaevolve, ). These agents go beyond code completion: they can generate entire programs or modules from natural language specifications, diagnose and fix bugs using compiler or test feedback, write and execute test cases, and refactor code for readability or performance. They can also invoke and interact with external tools such as compilers, debuggers, performance profilers, or version control systems, supporting an end-to-end development workflow.
@@
This emerging paradigm has the potential to fundamentally change how software is built and maintained. For example, an AI agent can take a natural language description of a feature and work through a series of steps, such as writing code, generating tests, running those tests, analyzing and fixing issues, and preparing a pull request. Some state-of-the-art coding agents have demonstrated the ability to continue working for hours while maintaining task consistency, avoiding deadlocks, and recovering from failed actions (qian2024chatdevcommunicativeagentssoftware, ; novikov2025alphaevolve, ). These systems can generate and test code, migrate software between frameworks, debug runtime failures, and integrate new features by decomposing complex goals into manageable subtasks (deng2024pentestgptllmempoweredautomaticpenetration, ; liu2023bolaa, ). This represents a clear shift from static, one-shot AI-based code generation to interactive, iterative, and tool-augmented workflows.
@@
Although progress has been fast, AI agentic programming is still in its early stages. Existing systems vary in architecture, autonomy, tool integration, and reasoning capabilities. There is no standard taxonomy, benchmark suite, or evaluation methodology. Moreover, multiple key challenges remain, including improving reliability (hou2024largelanguagemodelssoftware, ), reducing errors or hallucinations (gong2025language, ), handling tasks across different platforms and languages (DBLP:conf/pldi/polycoder22, ), and making sure these systems are safe and trustworthy in real use (yang2024swe, ).
@@
The success of AI coding agents depends heavily on their ability to interact effectively with external tools. However, today’s programming languages, compilers, and debuggers are fundamentally human-centric. They are not designed for automated, autonomous systems. These tools often abstract away internal states and decision-making processes to improve usability, ensure portability, and reduce cognitive load for human users (llvm, ; Maleki, ). While this abstraction benefits human developers, it may not fit AI agents, which require fine-grained, structured access to internal states, transformation sequences, and validation logic in order to reason about the effects of their actions (bi2024iterative, ). Without such access, AI agents struggle to diagnose failures, understand the implications of their changes, or recover from errors in a principled way. For instance, when a code transformation leads to a build failure, the agent needs more than just an error message - it must trace the failure to specific intermediate steps and understand why the changes caused the issue. Existing development environments do not provide hooks and feedback mechanisms to support this kind of iterative, tool-integrated reasoning.
Similarly, agentic coding systems benefit significantly from toolchains that support iterative development, state tracking, and rich feedback propagation - capabilities that most conventional tools do not expose. To operate effectively, AI agents may need access to internal compiler representations, transformation traces, symbolic information, and execution metadata. This raises a fundamental question: Are our current programming languages and software development tools still adequate in the era of AI agentic programming? Or is it time to rethink the design of programming languages, compilers, and debuggers to treat AI agents as first-class participants in the development process?
@@
These challenges show that AI agentic programming is not just a new way of using existing tools. It is a shift that exposes important gaps in how today’s software systems are designed. As the field evolves rapidly, there is a growing need to clarify its conceptual landscape, identify common patterns and system architectures, and assess the suitability of current development ecosystems. This is the right moment to step back, take stock of recent progress, and lay out the key questions that researchers and developers need to tackle next.
Therefore, this survey aims to provide a comprehensive overview of the emerging field of AI agentic programming. Specifically, it covers:
@@
  • A summary of current applications and practical use cases,
  • An analysis of evaluation strategies and benchmarking practices,
@@
We focus primarily on LLM-driven agentic systems for software development, though many insights extend to general task-oriented agents. Our goal is to chart the current landscape, clarify foundational concepts, and support the design of robust, efficient, and trustworthy AI agents for programming.
@@ 2.1. AI Agentic Programming
Figure 1. A representative workflow of an AI coding agent.
@@
AI agentic programming refers to a new programming paradigm in which LLM-based agents autonomously perform software development tasks. Unlike traditional code generation tools that produce outputs in a single step based on a static prompt (DBLP:conf/pldi/polycoder22, ), agentic systems operate in a goal-directed, multi-step manner. They reason about tasks, make decisions, use external tools (such as compilers, debuggers, and test runners), and iteratively refine their outputs based on feedback (DBLP:conf/nips/ShinnCGNY23, ; liu2024large, ; novikov2025alphaevolve, ). These agents can plan sequences of actions, adapt their strategies over time, and coordinate complex development workflows with limited or no human intervention.
@@
At its core, agentic programming combines the capabilities from natural language processing, external tool integration, and task planning. Figure 1 illustrates the architecture of a Copilot-style agentic programming system. At its core, the agent embeds an LLM within an execution loop, enabling interaction with the development environment. The agent receives natural language prompts from the user and gathers additional context from the operating system and the workspace (e.g., file summaries or environment state). This information is passed into the reasoning loop, where the LLM decomposes the task into subgoals, generates code or decisions, and determines whether to invoke external tools—such as reading/editing files or executing terminal commands. Tool outputs are returned to the loop and used as feedback for further refinement. This iterative process continues until the agent completes the task or reaches a stopping condition. Final results are streamed back to the user.
@@
AI agentic programming is characterized by several key properties. First, it emphasizes autonomy, where LLM-based agents can make decisions and take actions without continuous human supervision. Second, it is inherently interactive, as agents engage with external tools and environments during execution. Third, it supports iterative refinement, allowing agents to improve outputs based on intermediate feedback. Finally, it is goal-oriented, with agents pursuing high-level objectives (e.g., sub-tasks generated from the user inputs) rather than simply responding to one-shot prompts.
@@
Together, these features mark a departure from earlier forms of automation and code generation based on rules (xu2008rule, ), classical machine learning models (TraditionalCodeGenration, ) or one-shot LLM calling (DBLP:conf/pldi/polycoder22, ). AI agentic programming represents a change toward intelligent systems that actively participate in the software development process. This enables new capabilities in intelligent code assistance (github_copilot, ), autonomous debugging and testing (englhardt2024exploring, ; deng2024pentestgptllmempoweredautomaticpenetration, ), automated code maintenance (fan2023automated, ), and potentially even self-improving software systems (DBLP:journals/corr/abs-2405-15189, ; DBLP:journals/corr/abs-2310-02304, ).
@@ 2.1.1. Working example
As an example of AI agentic programming, consider a developer who tasks an AI coding agent with the following request: “Implement a REST API endpoint that returns the top 10 most frequently accessed URLs from a web server log file. Include unit tests and documentation.” This task requires integrating multiple software components, including file parsing, frequency analysis, web API implementation, testing, and documentation.
A high-level view of an agentic loop for solving this task is depicted in Figure 2. Here, an LLM begins by analyzing the natural language task and planning a sequence of actions. It first produces a Python function to parse the log file and count URL frequencies using a dictionary or a data analysis library like collections.Counter. Next, it implements a REST API endpoint using a web framework such as Flask, exposing a route like /top-urls that returns the computed result in JSON format. The LLM then writes unit tests and calls a Python interpreter to execute the generated Python script in the terminal and collects output to validate both the parsing logic and the API usage. It runs these tests using a tool like pytest, identifies failing cases, and refines the implementation. If a test fails due to a corner case (e.g., missing fields or malformed input), the LLM goes back to change the Python code, e.g., by adding input validation. It repeats this process - running tools, interpreting results, and modifying the code - until all tests pass. Finally, the LLM can generate documentation strings for each function and call an external tool like pdoc or Sphinx to produce human-readable API documentation. The process concludes when the agent validates that the tests pass, the API behaves as expected, and the documentation is complete.
This example illustrates the core features of agentic programming: autonomous planning, tool integration, iterative refinement, and goal-directed behavior. Unlike one-shot code generation, the agent interacts continuously with tools, learns from feedback, and adapts its actions to deliver a complete and functional software component.
@@ 2.2. Historical Context and Motivations
Figure 3. A diagram showing the evolution of coding agents from program synthesis, to code completion tools to AI coding agents
@@
As illustrated in Figure 3, the idea of automating software development has long been a goal in artificial intelligence and software engineering research (solar2008program, ). Early efforts in program synthesis aimed to generate correct-by-construction programs from formal specifications (gulwani2017program, ), while code completion tools sought to improve developer productivity by predicting likely code snippets based on contexts (github_copilot, ; tabnine, ; amazon-q-developer, ). These approaches, while impactful, typically relied on classical machine-learning models (TraditionalCodeGenration, ), handcrafted ruless (xu2008rule, ), or statistical techniques with limited generalizations (bruch2009learning, ).
@@
The advent of large-scale pre-trained LLMs such as Codexs (openai2021codex, ), GPTs (openai2024gpt4o, ), Claudes (claude2024, ), LLaMas (meta2024llama3, ) and StarCoder (lozhkov2024starcoder, ) marked a major turning point. These models demonstrated strong zero-shot and few-shot capabilities in generating code (DBLP:conf/nips/Le0GSH22, ; peng2024perfcodegen, ), translating between programming languages (DBLP:conf/emnlp/2020pymt5, ; DBLP:conf/mipro/GuzziKPGB24, ), and answering complex programming questions with little or no task-specific fine-tuning (DBLP:journals/corr/abs-2203-07814, ; coignion2024performance, ). Their ability to understand and generate natural language and code made them a good fit for software development tasks beyond basic code completion (husein2025large, ), including documentation generation (dvivedi2024comparative, ), test synthesis (jorgensen2024large, ), and bug detection and fixing (englhardt2024exploring, ; deng2024pentestgptllmempoweredautomaticpenetration, ; kang2025explainable, ).
As these models became more capable, a new opportunity emerged: using LLMs not just as passive code generators, but as autonomous agents that could reason about goals, invoke tools, and refine their outputs over multiple steps. This shift leads to the paradigm of AI agentic programming, where models operate as task-driven entities capable of planning, interacting with compilers and debuggers, and self-correcting based on feedback.
@@
Several trends motivated this change. First, real-world software development often requires iterative problem solving, tool use, and adaptation, which single-step code generation cannot handle effectively (liu2024large, ; li2024survey, ). Second, the rise of prompt engineering and structured prompting techniques (e.g., ReAct, chain-of-thought, scratchpads) enabled LLMs to reason more effectively over multiple steps (yao2023react, ; guo2025deepseek, ). Third, the increasing availability of APIs, command-line tools, and language server protocols made it possible to integrate LLMs into full-stack development environments (anthropic_claude_cli, ; google_cloud_cli, ; anthropic_mcp_spec, ).
These developments prompted a rethinking of how LLMs could be deployed-not just as smart coding assistants (github_copilot, ; tabnine, ), but as semi-autonomous agents capable of carrying out software engineering tasks with minimal human supervision. The resulting systems promise to augment developer productivity, lower the barrier to entry for software creation, and open new avenues for automation in programming workflows.

### v2

@@ Abstract.
AI agentic programming is an emerging paradigm where large language model (LLM)-based coding agents autonomously plan, execute, and interact with tools such as compilers, debuggers, and version control systems. Unlike conventional code generation, these agents decompose goals, coordinate multi-step processes, and adapt based on feedback, reshaping software development practices. This survey provides a timely review of the field, introducing a taxonomy of agent behaviors and system architectures and examining relevant techniques for planning, context management, tool integration, execution monitoring, and benchmarking datasets. We highlight challenges of this fast-moving field and discuss opportunities for building reliable, transparent, and collaborative coding agents.
@@ 1. Introduction
The software development paradigm is changing rapidly with the rise of large language models (LLMs) (Hou et al., 2024). These models enable artificial intelligence (AI) systems that not only translate natural language descriptions into code snippets (Jiang et al., 2025) but also understand task requirements, interact with development tools, and iteratively refine their outputs to produce complex software (Alshahwan et al., 2024; Qian et al., 2024). Recent studies suggest that developers now use LLMs routinely to assist in daily coding tasks (Hou et al., 2024; Fan et al., 2023b; Gong et al., 2025). Unlike traditional code generation tools (Allamanis et al., 2018a) that respond to a single prompt with a static code snippet, emerging AI coding agents are designed to operate within dynamic software environments, performing iterative, tool-augmented tasks to achieve complex goals.
@@
This shift has given rise to a new programming paradigm, AI agentic programming, where LLM-based coding agents can autonomously plan, execute, and refine software development tasks (Gao et al., 2023; Novikov et al., 2025). Unlike conventional code-completion tools (Int, 2025; LSP, 2025; Bruch et al., 2009a), which primarily assist with local suggestions, these agents are capable of generating entire programs or modules from natural language specifications, diagnosing and fixing bugs using compiler or test feedback, writing and executing test cases, and refactoring code for readability or performance. Beyond code generation, they can also orchestrate external tools, such as compilers, debuggers, performance profilers, and version control systems, supporting an end-to-end development workflow.
@@
This emerging paradigm has the potential to fundamentally change how software is built and maintained. For example, an AI agent can take a natural language description of a feature and work through a series of steps, such as writing code, generating tests, running those tests, analyzing and fixing issues, and preparing a pull request (Qian et al., 2024). Some state-of-the-art coding agents, like Anthropic’s Calude Opus 4 (Cal, 2025), have demonstrated the ability to continue working for hours while maintaining task consistency, avoiding deadlocks, and recovering from failed actions (Qian et al., 2024; Novikov et al., 2025). These systems can generate and test code, migrate software between frameworks, debug runtime failures, and integrate new features by decomposing complex goals into manageable subtasks (Deng et al., 2024; Liu et al., 2023d). This represents a clear shift from static, one-shot AI-based code generation to interactive, iterative, and tool-augmented workflows.
@@
Although progress has been fast, AI agentic programming is still in its early stages. Existing systems vary in architecture, autonomy, tool integration, and reasoning capabilities. There is no standard taxonomy, benchmark suite, or evaluation methodology. At the same time, multiple key challenges remain. These include ensuring reliability and robustness in dynamic environments (Hou et al., 2024), mitigating errors and hallucinations in generated code (Gong et al., 2025), extending support beyond dominant languages such as Python to diverse platforms and software ecosystems (Xu et al., 2022), and embedding safety, trust, and accountability into autonomous behaviours (Yang et al., 2024b).
@@
The success of AI coding agents also depends heavily on their ability to interact effectively with external tools. However, today’s programming languages, compilers, and debuggers are fundamentally human-centric (Marron, 2023; Aho et al., 2006). They are not designed for automated, autonomous systems. These tools often abstract away internal states and decision-making processes to improve usability, ensure portability, and reduce cognitive load for human users (Lattner and Adve, 2004a; Maleki et al., 2011b). While this abstraction benefits human developers, it may not fit AI agents, which require fine-grained, structured access to internal states, transformation sequences, and validation logic in order to reason about the effects of their actions (Bi et al., 2024). Without such access, AI agents struggle to diagnose failures, understand the implications of their changes, or recover from errors in a principled way. For instance, when a code transformation leads to a build failure, the agent needs more than just an error message - it must trace the failure to specific intermediate steps and understand why certain code edits or actions caused the issue. Existing development environments do not provide hooks and feedback mechanisms to support this kind of iterative, tool-integrated reasoning. Similarly, agentic coding systems would benefit from toolchains that support iterative development, state tracking, and rich feedback propagation - capabilities that most conventional tools do not expose. To operate effectively, AI agents may need access to internal compiler representations, transformation traces, symbolic information, and execution metadata.
@@
These challenges show that AI agentic programming is not just a new way of using existing tools. It is a shift that exposes important gaps in how today’s systems software is designed. As the field evolves rapidly, there is an urgent need to clarify its conceptual landscape, identify common patterns and system architectures, and assess the suitability of current development ecosystems. This is the right moment to step back, take stock of recent progress, and lay out the key questions that researchers and developers need to tackle next. Therefore, this survey aims to provide a comprehensive overview of the emerging field of AI agentic programming. Specifically, it covers:
@@
  • A summary of the behavior dimensions, operating modes, evaluation strategies and benchmarking practices of AI coding agents,
@@
We focus primarily on LLM-driven agentic systems for software development, though many insights extend to general AI agents in other domains like information retrieval (Li et al., 2025a). Our goal is to chart the current landscape, clarify foundational concepts, and support the design of robust, efficient, and trustworthy AI agents for programming.
@@ 2.1. AI Agentic Programming
Figure 1. An example workflow of an AI coding agent.
@@
AI agentic programming refers to a new programming paradigm in which LLM-based agents autonomously perform software development tasks. Unlike traditional code generation tools that produce outputs in a single step based on a static prompt (Xu et al., 2022), agentic systems operate in a goal-directed, multi-step manner. They reason about tasks, make decisions, use external tools (such as compilers, debuggers, and test runners), and iteratively refine their outputs based on feedback (Shinn et al., 2023; Liu et al., 2024e; Novikov et al., 2025). These agents can plan sequences of actions, adapt their strategies over time, and coordinate complex development workflows with limited or no human intervention.
@@
At its core, agentic programming combines the capabilities of natural language processing, external tool integration, and task planning. Figure 1 illustrates the architecture of a GitHub Copilot-style agentic programming system (GitHub, 2021). At its core, the agent embeds an LLM within an execution loop, enabling interaction with the development environment. The LLM receives natural language prompts from the user and gathers additional context from the operating system and the workspace (e.g., file summaries or environment state). This information is passed into the reasoning loop, where the LLM decomposes the task into subgoals, generates code or decisions, and determines whether to invoke external tools like reading/editing files or executing terminal commands. Tool outputs are returned to the loop and used as feedback for further refinement. This iterative process continues until the agent completes the task or reaches a stopping condition. Final results are streamed back to the user.
@@
A typical AI agentic programming system is characterized by several key properties. First, it emphasizes autonomy, where LLM-based agents can make decisions and take actions without continuous human supervision. Second, it is inherently interactive, as agents engage with external tools and environments during execution. Third, it supports iterative refinement, allowing agents to improve outputs based on intermediate feedback. Importantly, it is goal-oriented, with agents pursuing high-level objectives (e.g., sub-tasks generated from the user inputs) rather than simply responding to one-shot prompts.
@@
Together, these features mark a departure from earlier forms of automation and code generation based on rules (Xu, 2008), classical machine learning models (Allamanis et al., 2018a) or one-shot LLM calling (Xu et al., 2022). AI agentic programming represents a change toward intelligent systems that actively participate in the software development process. This enables new capabilities in intelligent code assistance (GitHub, 2021), autonomous debugging and testing (Englhardt et al., 2024; Deng et al., 2024), automated code maintenance (Fan et al., 2023a), and potentially even self-improving software systems (Huang et al., 2024b; Zelikman et al., 2024).
@@ 2.1.1. Working example
As an example of AI agentic programming, consider a developer who tasks an AI coding agent with the following request: “Implement a REST API endpoint that returns the top 10 most frequently accessed URLs from a web server log file. Include unit tests and documentation.” This task requires integrating multiple software components, including file parsing, frequency analysis, web API implementation, testing, and documentation. A high-level view of an agentic loop for solving this task is depicted in Figure 2. Here, an LLM begins by analyzing the natural language task and planning a sequence of actions. It first produces a Python function to parse the log file and count URL frequencies using a dictionary or a data analysis library like collections.Counter. Next, it implements a REST API endpoint using a web framework such as Flask, exposing a route like /top-urls that returns the computed result in JSON format. The LLM then writes unit tests and calls a Python interpreter to execute the generated Python script in the terminal and collects output to validate both the parsing logic and the API usage. It runs these tests using a tool like pytest, identifies failing cases, and refines the implementation. If a test fails due to a corner case (e.g., missing fields or malformed input), the LLM goes back to change the Python code, e.g., by adding input validation. It repeats this process - running tools, interpreting results, and modifying the code - until all tests pass. Finally, the LLM can generate documentation strings for each function and call an external tool like pdoc or Sphinx to produce human-readable API documentation. The process concludes when the agent validates that the tests pass, the API behaves as expected, and the documentation is complete. This example illustrates the core features of agentic programming: autonomous planning, tool integration, iterative refinement, and goal-directed behavior. Unlike one-shot code generation, the agent interacts continuously with tools, learns from feedback, and adapts its actions to deliver a complete and functional software component.
@@ 2.2. Historical Context and Motivations
Figure 3. The evolution of coding agents from program synthesis, to code completion tools, to AI coding agents.
Figure 4. Evolution of code generation approaches.
2.2.1. A systems perspective
As illustrated in Figure 3, the idea of automating software development has long been a goal in artificial intelligence and software engineering research (Solar-Lezama, 2008). In some respects, the rise of AI agents echoes the historical evolution of operating systems. Early operating systems introduced processes as the basic active execution units, capable of running independently. These evolved into threads, lightweight abstractions that enabled efficient concurrency and finer-grained scheduling (Wollheim, 1984). Later, with object-oriented design (Riel, 1996), objects became the dominant unit of modularity - encapsulating state and behaviour, but remaining fundamentally passive. Active objects (Nierstrasz, 1987, 1993) extended the object model by incorporating their own thread of control and asynchronous message passing, allowing them to act autonomously rather than merely responding to external calls—an early precursor to modern agents.
LLM-based AI agents can be seen as the next step in this trajectory. Like processes and threads, they are active entities, but unlike objects, they do not wait passively for instructions at every step. Instead, they can plan, act, and coordinate tasks proactively across diverse contexts. Their behavior is not restricted to fixed, handwritten policies but is driven by LLMs, enabling them to adapt dynamically, invoke external tools, and collaborate with both humans and other agents (Li et al., 2024e). In this sense, agentic programming represents a new stage in the long-standing pursuit of making software development more autonomous, where code generation, debugging, and optimization are no longer fully hand-crafted but emerge through iterative interactions between humans, tools, and intelligent agents.
2.2.2. Algorithm evolution of agents
@@
Figure 4 shows the evolution of algorithms. Early efforts in program synthesis aimed to generate correct-by-construction programs from formal specifications (Gulwani et al., 2017), while code completion tools sought to improve developer productivity by predicting likely code snippets based on contexts (GitHub, 2021; Tabnine, 2025; Amazon Web Services, 2025). These approaches, while impactful, typically relied on classical machine-learning models (Allamanis et al., 2018a), handcrafted ruless (Xu, 2008), or statistical techniques with limited generalizations (Bruch et al., 2009b).
@@
The advent of large-scale pre-trained LLMs such as Codexs (OpenAI, 2021), StarCoder (Lozhkov et al., 2024), LLaMas (Meta, 2024), GPTs (OpenAI, 2024b), Claudes (Anthropic, 2024b), Gemini (et al., 2025), Qwen (Hui et al., 2024), and Deepseek (Guo et al., 2025), marked a major turning point. These models demonstrated strong zero-shot and few-shot capabilities in generating code (Le et al., 2022; Peng et al., 2024a), translating between programming languages (Clement et al., 2020; Guzzi et al., 2024), and answering complex programming questions with little or no task-specific fine-tuning (Li et al., 2022; Coignion et al., 2024). Their ability to understand and generate natural language and code made them a good fit for software development tasks beyond basic code completion (Husein et al., 2025), including documentation generation (Dvivedi et al., 2024), test synthesis (Jorgensen et al., 2024), and bug detection and fixing (Englhardt et al., 2024; Deng et al., 2024; Kang et al., 2025). As these models became more capable, a new opportunity emerged: using LLMs not just as passive code generators, but as autonomous agents that could reason about goals, invoke tools, and refine their outputs over multiple steps. This shift leads to the paradigm of AI agentic programming, where models operate as task-driven entities capable of planning, interacting with compilers and debuggers, and self-correcting based on feedback.
@@
Several trends motivated this change. First, real-world software development often requires iterative problem solving, tool use, and adaptation, which single-step code generation cannot handle effectively (Liu et al., 2024e; Li et al., 2024e). Second, the rise of prompt engineering and structured prompting techniques (e.g., ReAct, chain-of-thought, scratchpads) enabled LLMs to reason more effectively over multiple steps (Yao et al., 2023c; Guo et al., 2025). Third, the increasing availability of APIs, command-line tools, and language server protocols made it possible to integrate LLMs into full-stack development environments (Anthropic, 2025a; Cloud, 2025b; Anthropic, 2025c). These developments prompted a rethinking of how LLMs could be deployed-not just as smart coding assistants (GitHub, 2021; Tabnine, 2025), but as semi-autonomous agents capable of carrying out software engineering tasks with minimal human supervision.